Tamika has $2.50 in quarters in her pocket,including four stgate quarters. She reaches into her pocket and takes out two quarters, one at a time. What is the probablility that they are both state quarters?

Answer by stanbon(75887)

Tamika has $2.50 in quarters in her pocket,including four state quarters. She reaches into her pocket and takes out two quarters, one at a time. What is the probablility that they are both state quarters?
----
She has 10 quarters; 4 of them are state quarters.
----
# of ways to pick 2 state quarters: 4C2 = 6
# of ways to pick 2 quarters randomly: 10C2 = 45
--------------------------------------------------
P(2state quarters) = 6/45 = 2/15
====================================
Cheers,
Stan H.
=============